OpenStudy (anonymous):

Which ordered pair is in Quadrant IV? A. (–3, –5) B. (–3, 5) C. (5, –3) D. (5, 3)

OpenStudy (anonymous):

Ah! i hate when im on a rush

OpenStudy (anonymous):

Quadrant 4 is the box at the bottom right

OpenStudy (anonymous):

O

OpenStudy (anonymous):

So which one is at the bottom right box? https://www.desmos.com/calculator/mcoc3impz5

OpenStudy (anonymous):

@PlanSlam

OpenStudy (anonymous):

@PlanSlam Did you get your answer yet?

OpenStudy (conqueror):

@PlanSlam We cannot give you the answer as it is against the code of conduct. And you should not be looking for the answer either.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the distance between points (9, 4) and (–3, 4) on a coordinate plane?

OpenStudy (anonymous):

what do i do for this one?

OpenStudy (anonymous):

do i add 9 +-3?

OpenStudy (igreen):

Use the distance formula.. \(d = \sqrt{(x_2 - x_1)^2 + (y_2-y_1)^2)}\) Plug in the points: \(d = \sqrt{(-3 - 9)^2 + (4-4)^2)}\) Simplify

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

will it be 6,4?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Actually it should be one number..

OpenStudy (igreen):

What's -3 - 9?

OpenStudy (igreen):

@PlanSlam